OpenStudy (anonymous):

.A wave with a frequency of 0.5 Hz and a speed of 10 m/s has a wavelength of

OpenStudy (anonymous):

L=U/u=10/0,5=20

OpenStudy (anonymous):

20m

OpenStudy (anonymous):

well speed = frequency * wavelength => wavelength = speed / frequency and here speed = 10 m/s and frequency = 0.5 Hz => wavelength = 10/0.5 = 100/5 = 20 and since SI unit of wavelength is meter therefore answer = 20 m